### v2

# Towards Post-mortem Data Management Principles for Generative AI

Abstract
Foundation models, large language models (LLMs), and agentic AI systems all rely on vast corpora of user data. The use of such data for training these models has raised ongoing concerns around data ownership, copyright issues and potential harm may caused. In this work, we explore a related but less examined angle: the ownership rights of data belonging to deceased individuals. To do so, we first examine the current landscape of data management and privacy rights for deceased individuals as defined by privacy policies of major technology companies, and by privacy regulations such as the EU AI Act. Based on our analysis, we propose three post-mortem data management principles to guide the protection of deceased individuals’ data rights. Finally, we discuss potential future work and recommendations for policymakers and privacy practitioners to deploy these principles, alongside technological solutions to implement and audit them in practice.
1 Introduction
The latest advances in generative AI (Gen-AI) have produced powerful large language model (LLM) systems and agentic AI platforms, including ChatGPT OpenAI (2025), Google Gemini Google (2025), Anthropic Claude Anthropic (2025), Microsoft Copilot Microsoft (2025), Meta LLaMA Meta (2025), Replika AI Luka Inc. (2025), and Character.AI Character.AI (2021), among others. They can generate and interpret text, images, video, audio, and can even imitate a person’s voice, image, and creative style. Gen-AI systems depend heavily on large volumes of user-generated data, including personal information and  Carlini et al. (2023); Nasr et al. (2025); OpenAI (2023); Touvron et al. (2023); Zhang et al. (2023). The use of such data raises ongoing concerns about data ownership, copyright, privacy, and potential harms  Croitoru et al. (2024); Wang et al. (2023); Yu et al. (2025); Nie et al. (2025); Nuñez (2025).
Under existing privacy regulations, for example under the European Union’s General Data Protection Regulation (GDPR) European Union (2016), the California Consumer Privacy Act (CCPA) California State Legislature (2018), or the Brazilian Lei Geral de Proteção de Dados (LGPD) Government (2018), users have the right to file complaints with regulatory bodies, pursue legal action, request the deletion of their personal data, and opt out of certain types of data processing to safeguard their digital privacy. However, protections for deceased users are limited. Privacy regulations generally do not apply to deceased users, thus their digital data often remains accessible, risking privacy violations, unauthorized use, and even non-consented digital cloning that may cause harm to their legacy and memory, and harm their surviving family and friends Morris and Brubaker (2025).
A notable example is Character.AI case Farokhmanesh and Goode (2024), where Jennifer, a young woman who was killed, whose digital presence was used without consent to create an AI chatbot impersonating her. While policy violations were acknowledged by Character.AI following public outcry, such incidents underscore the urgent need for clearer protections for deceased individuals’ privacy rights. Research highlights Malas (2024) a growing interest in “deadbots” or “griefbots,” AI chatbots that replicate deceased person’s language and personality from social media data. While potentially comforting at first, they may later become burdensome; e.g., by sending unwanted messages that make survivors feel “stalked by the dead.”
Figure 1: Overview of our approach : (i) analysis phase, examining current regulations, industry practices, and research to draw insights; (ii) based on this analysis, we propose three principles; and (iii) implementation phase where we provide recommendations on implementing the principles through regulatory and technological solutions
Recently Meta is compensating individuals to record facial expressions, speech, and small talk for building realistic virtual avatars under “Project Warhol”  Web (2025). While participation is voluntary and consent-based, this practice can be concerning regarding our problem space. Once created, these digital avatars may persist indefinitely, extending a person’s digital footprint beyond their lifetime. Without clear protections, "realistic avatars" of deceased individuals can be exploited without consent, may result in privacy violations, identity misuse, unauthorized commercial exploitation, and manipulation as well as reputational harm and psychological distress upon surviving relatives.
Contributions. To address those issues, we investigate an underexplored aspect of data privacy: the data management rights of deceased individual’s data in the context of Gen-AI systems. Our work makes following contributions:
  • We highlight the overlooked issue of deceased individuals’ data rights in Gen-AI systems.
  • We systematically examine current regulations and industry practices to identify gaps in post-mortem data protection in the age of Gen-AI.
  • We propose three Post-mortem data management principles to manage the collection, retention, and use of post-mortem digital data in AI systems, ensuring ethical handling and preventing misuse. We also recommend regulatory and technical solutions to deploy them.
2 Current Landscape and Insights on Post-Mortem Data Management
Privacy and AI Regulations. The EU AI Act European Parliament (2025) imposes transparency obligations on Gen-AI systems, even when they are not classified as high-risk. These include disclosing that content is AI-generated, implementing safeguards to prevent illegal content generation, and publishing summaries of copyrighted material used during training. More advanced general-purpose models, like GPT-4 which may pose systemic risks, must undergo rigorous evaluations.Similarly, starting January 1, 2026, California’s AI Transparency Act (SB 942) requires Gen-AI systems producing multimedia content, for example, synthetic voices, deepfakes, or images, with over one million monthly users in California to include disclosures and detection tools. The law mandates visible labels, embedded watermarks, and free verification tools, with enforcement by the California Attorney General Partow-Navid (2025). Recent U.S. AI Action Plan also emphasizes improving AI accessibility and robustness within federal agencies. It calls for expanding secure access to federal data, developing a government wide AI procurement toolbox, and promoting AI systems that are resilient to adversarial attacks. However, it does not address data ownership or copyright concerns in training Gen-AI systems  U.S. Government (2025).
Recent privacy laws regulate the data subjects’ right to be forgotten as per GDPR Mantelero (2013) or right to delete/erase as per CCPA California State Legislature (2018) or LGPD Government (2018), stating that a data subject has the right to request the deletion of personal information collected by a company. Further, for machine learning models, the UK Information Commissioner’s Office (ICO) issued draft guidance that stated that a data deletion request may require retraining or deleting the model entirely. However, most privacy and AI laws do not apply to data from deceased individuals. For instance, the EU AI act defines deepfakes as "AI-generated or manipulated image, audio or video content that resembles existing persons … " European Parliament (2025), thus not considering imitations of deceased individuals as deepfakes.
Industry Practices. Current industry handles post-mortem data mainly at the account level (see Appendix B). To analyze data management policy for AI agents, we selected from Table 1 (Appendix D) based on popularity Topics (2025); Carolan et al. (2025) and relevance to post-mortem data may use/sustain, including agents from the General AI Assistant, AI Companion, and Mental Health categories that process personal data and capture user behavior or personality. We review these agents data practices by examining their privacy policies and how they handle user data.
Major AI platforms vary in how they handle user data post-interaction. For instance, OpenAI allows users to access, manage, or delete their data OpenAI (2024). Chat history is generally removed within 30 days, and while user content is by default used for model improvement, users can opt out through data controls. Similarly, Anthropic retains Claude prompts and outputs for 30 days post-deletion for safety and compliance, unless legally required otherwise. Enterprise users can define custom retention and deletion policies via organizational controls Inc. (2024). Character.AI Character.AI (2021), allows creating chatbots based on fictional or real individuals, retains user conversations for personalization, model improvement, and moderation. Users can request deletion of chat history, though some data may persist for legal or safety with no clear retention timeline Character.AI (2025). The platform enforces Digital Millennium Copyright Act (DMCA) DCM (1998) take downs for copyright-infringing characters, but enforcement is often slow and unauthorized bots persist, highlight the need for stronger protections of deceased individuals’ data and safeguards against misuse. Replika AI Luka Inc. (2025) provides personalized companionship through adaptive conversations, role-play, and customizable personalities, enhanced by voice and AR features. User data, including conversational messages, is used to train models that are prominently featured in Replika’s promotional materials. While not directly shared with advertisers, this practice raises ethical concerns, as the platform targets its advertising towards vulnerable or lonely users Piispanen et al. (2024). Coupled with the digital footprints of deceased individuals, these practices could be used to attract relatives, friends, or close contacts of the deceased to engage with the platform which is concerning.
Mental health agents such as Woebot Woebot Health (2025) and Finch Finch Care (2025) adopt different approaches. Woebot allows users to access, correct, or delete their data and follows HIPAA-aligned security practices, but it is not legally bound by HIPAA unless offered through a covered healthcare provider, leaving limited legal privacy protections Health (2022). Finch permits users to request deletion and exercise rights of access and correction, while retaining the ability to use aggregated, non-identifiable data for analytics, marketing, and service improvement Finch (2025).
Related Research. Prior works have examined post-mortem privacy, highlighting the lack of legal control over deceased individuals’ data and the ethical concerns. Studies have documented user preferences for managing digital legacies, including data deletion, commercial use, and post-mortem AI agents, while emphasizing potential harms to survivors. More details are provided in Appendix A.
Analytical Insights. Despite their remarkable capabilities, Gen-AI models pose significant privacy and ethical risks. As noted in Sections 1 and Appendix A, they have been used to replicate deceased individuals, illustrating AI afterlife and generative ghost concepts. Existing regulations, including GDPR, CCPA and EU AI Act European Parliament (2025), protect living users by enabling data deletion and control over AI model influence, but largely exclude deceased individuals and leaving it unregulated. While current privacy laws focus on transparency, safety, user rights, content disclosure, harm prevention, and watermarking to prevent misinformation, they provide no explicit protections for post-mortem data rights. Persistent digital traces of deceased individuals can be misused for misinformation, identity replication, or monetization, raising urgent ethical and legal concerns. Recalling industry practices, major companies provide users with options to manage their data after death. However, these options are limited to social media accounts and do not extend to AI platforms or MLaaS services. Major AI platforms provide users with some control over data deletion. However, these options are designed for living users and do not address post-mortem data management.
3 Proposed Principles
Based on Section 2, we propose the following 3 principles for post-mortem data in Gen-AI systems:
1. Right to be forgotten or data deletion. As highlighted in section 2, individuals may wish for their data to be deleted after death. Mechanisms should ensure both the deletion of personal data and the removal of its influence from AI models, providing users with options to specify post-mortem deletion or deletion after a defined period of inactivity. After verifying users’ death, (preferably by a designated legacy contact selected by the user during their lifetime), service provider must initiate both conventional data deletion and the removal of the deceased’s influence on data retrieval.
2. Data Inheritance and Ownership. Individuals may choose not to have their data deleted after death (see appendix A) if it can be monetized or if they wish to preserve an AI afterlife for their loved ones, opting instead to pass their data rights to their heirs. However, as personal data may contain sensitive information, inheritance may proceed through three approaches: deletion on behalf, transferring the actual data to heirs, or providing heirs with the monetized value of the data without direct access.
3. Purpose Limits & Harm Prevention. In cases where deceased individuals consent to donate their data for research, societal benefit, or other public purposes, explicit agreements should be established to ensure compliance with legal and ethical standards. In such instances, clear agreements should be established, incorporating the following criteria: (i) transparency regarding the intended use of the data, (ii) purpose limitations that uphold the dignity and post-mortem rights of the deceased, (iii) safeguards to prevent harm to the deceased’s legacy and to surviving relatives, and (iv) privacy protections. Organizations managing donated data should implement binding agreements that clearly specify permissible uses, aligned with consent provided by the deceased and relevant regulations.
4 Discussion
Regulatory Enforcements. We recommend increased regulations and adding protections for deceased individuals’ data, following our proposed principles. AI agents’ privacy policies should be required to disclose how personal data is processed after death and clearly state individuals’ rights for post-mortem data management. Concerning data deletion requests by heirs, we recommend a standardized process: within 30 days (matching the GDPR response period) the provider must delete all associated data, e.g., search indexes, and, within similar timeline, remove personal data influence from AI models and send a compliance report to the legacy contact.
Additionally, clear guidelines should be established regarding data ownership and digital assets inheritance, with a focus on security and privacy. To support this, we suggest a streamlined digital will mechanism that empowers individuals to determine the post-mortem management of their digital assets. Much like testamentary freedom in estate law Goralewicz (2024), this mechanism would allow a person to exercise control over whether their data should be deleted, retained, shared, or monetized after death.
In the case of data donations, regulations should prohibit use for targeted advertising, political persuasion, or voice/image cloning that may misrepresent the deceased or cause harm, reflecting emerging AI regulations such as the EU AI Act on synthetic content European Parliament (2025). Institutional review boards or human oversight committees should verify compliance before any publication, sharing, or AI model training to uphold both ethical and legal responsibilities. See Appendix C for further discussion on continuation of post-mortem data use, to comply with the 2nd and 3rd principles.
Technical Solutions. Our proposed post-mortem principles can be implemented through two main technical approaches: regular system audits and the integration of privacy and safety by design. To enforce the first principle, data deletion requests must be honored upon user initiation. When user data is used for training or fine-tuning AI models, deletion must extend beyond the removal of raw data to include machine unlearning Hu et al. (2024); Bourtoule et al. (2021), ensuring that the user’s data no longer influences model behavior. Furthermore, rigorous evaluation mechanisms, such as 3rd party model audits Shi et al. (2025), should be implemented to verify that data deletion requests have been fully and effectively executed.
Digital will solutions help fulfill our second principle. For instance, Chen et al. (2025) introduce a digital will based on attribute based encryption, protecting user data privacy and offering fine-tuning of data sharing. To enable monetization of deceased users’ data without direct access, standardized approaches may be employed, such as pricing models, metadata, product taxonomies, and regression analyses for valuation Azcoitia et al. (2023), administered by a trusted third party, in combination with cryptographic methods Catalano (2014) that preserve ownership and identifiability without revealing underlying content.
To enforce our third principle, privacy and safety by design approaches should be adopted. Techniques such as data minimization, anonymization, or differential privacy (DP) Dwork and Roth (2014) can help privacy preserving data sharing and computing. Watermarking İşler et al. (2024) or canary injection Meeus et al. (2025) can be applied to donated post-mortem datasets, with watermarks indicating the copyright holder or distributing organization. Additionally, technical safeguards and tools Chi et al. (2024); Inan et al. (2023), such as content classification and prompt/response monitoring, can be adapted to detect and reduce harmful or disrespectful outputs, including content that may cause personalized harm to surviving relatives.
Limitations and Future Work. Our analysis is limited to a subset of AI agents, LLMs, and privacy laws, and does not cover all jurisdictions or platforms. We did not empirically evaluate technical solutions for enforcing the proposed principles. Future work includes auditing Gen-AI systems to measure: (i) the extent of sensitive data memorization, (ii) the effectiveness of current unlearning methods, and (iii) the potential for harm from Gen-AI using post-mortem data. Additionally, exploring efficient methods to deploy these principles constitutes a direction for future work.
Acknowledgments and Disclosure of Funding
We would like to thank Jeremy Epstein (Co-Director, ICARIS Center, Georgia Tech and former AD, White House, OSTP) for his valuable insights and discussion, particularly on the continuation of post-mortem data use. We would also like to thank Devris Isler (IMDEA Software Institute) for his initial discussions on data inheritance.
Appendix A Related Work
Legal Scholarship. Establishing regulations on deceased individuals’ data has been discussed by several authors Ashley (2020); Buitelaar (2017); Harbinja (2017); Allen and Rothman (2024), specifically concerning privacy rights. Authors argue that although people generate and acquire large volumes of digital assets, including personal and sensitive data, they are given no control over how this data is handled after they die. Allen and Rothman (2024) define post-mortem privacy as the protection of an individual’s privacy interests after death, focusing on the control of their personal data, image, voice, and other identifying attributes. Harbinja (2017) advances that establishing post-mortem privacy logically stems from the recognition of one’s autonomy. In addition, personal data can still affect the post-mortem portrayal of the deceased and emotional wellbeing of surviving family members. Allen and Rothman (2024) argue that, in the absence of law that protects post-mortem privacy, control over such data is left to the discretion of technology platforms that vary across platforms and may mismanage the data.
User studies. Individuals’ thoughts and preferences on post-mortem data management options have been recorded in multiple studies Chen et al. (2021); Pfister (2017); Lira et al. (2023); Reeves et al. (2024); Holt et al. (2021); Snow and Barry (2024); Nakagawa and Orita (2024). Several authors examined people’s opinion on different designs for digital legacy preparation and sharing Chen et al. (2021); Pfister (2017); Holt et al. (2021). It was found that most people start thinking about preparing a digital legacy after an emotional trigger, for instance the death of a loved one. A larger survey on 1020 Australian participants found that most do want some amount of control over how their data will be managed and shared after their death Reeves et al. (2024).
Lira et al. Lira et al. (2023) report on the perspective of heirs, and interview individuals who inherited digital assets. They focus on transferring social media account ownership or memorializing a social media page, for instance on Facebook. Two surveys seeked reactions for different outcomes of data after death: data donation Snow and Barry (2024) and the creation of immortal digital personalities Nakagawa and Orita (2024). Snow et al. propose the concept of a "data donor card", and surveyed 165 people in Dublin in 2023. They introduce discussion points, and highlight a large gap between legislation and opinion: even though most participants considered that data remains "personal data" even after one’s death, most legislation does not apply to data originating from deceased individuals. Nakagawa et al. Nakagawa and Orita (2024) explore the commercialization of “immortal digital personalities,” where deceased individuals’ personal data is used by Gen-AI systems to create interactive digital representations of the deceased individuals. Through a large-scale survey involving 2749 Japanese participants, the study reveals diverse post-mortem preferences: while many preferred deletion of their data, notably only about 20% of respondents indicated they would allow commercial use of their digital footprint if compensated during their lifetime. They emphasize that managing such legacy data is vital to uphold the deceased’s dignity and intentions, and identify risks of commodifying these identities as post-mortem entertainment
[…]
Prior works Morris and Brubaker (2025); Lei et al. (2025); Hollanek and Nowaczyk-Basińska (2024) explore users’ expectations, and ethical concerns regarding AI-generated post-mortem agents. Lei et al. Lei et al. (2025) explore AI afterlife, an emerging and dynamic form of digital legacy that leverages Gen-AI and differs from traditional static digital legacy. They emphasize maintaining identity consistency of the deceased while examining how user attitudes are shaped by personal, familial, technological, and social factors. Morris et al. Morris and Brubaker (2025) investigate generative ghosts, AI agents capable of producing novel content rather than merely replicating the deceased, and identify several risks, including mental health issues (complicated grief, anthropomorphism, “second deaths”), reputational harms (privacy breaches, hallucinated content), security threats (identity theft, hijacking, malicious ghosts), and socio-cultural impacts (changes to relationships, labor, and religious practices).
Appendix B Industry Practices to Manage Post-mortem Data
Google offers options to delete a deceased user’s account, transfer any associated funds, or request access to their data, provided that a valid request is made by a close relative or authorized representative Google (2024). Apple’s Digital Legacy program Apple (2024) allows designated contacts to access selected iCloud data with an access key and death certificate. However, designated contacts cannot learn Keychain passwords or payment data of the deceased person Apple (2024). Microsoft requires a court order or subpoena to access a deceased user’s account from Outlook, One-Drive, or other account data owned by Microsoft Microsoft (2024). X will deactivate an account when a verified family member or estate provides documentation but will not allow access to the contents of the account X (formerly-Twitter). Meta lets users assign legacy contacts to manage memorialized accounts Meta (2024). Legacy contacts can post a message, handle friend requests, update photos, download shared content (if allowed), and delete the account. They cannot log in, edit past posts, read messages, or remove friends. Overall, most companies focus on basic account management after a users’ death such as memorializing the profile, deleting their account, or allowing families to download some of their data.
Appendix C Discussion on Continuation of Post-Mortem Data Use
The continued use of post-mortem data, supported by our second and third principles, raises a key issue: the duration of such use. One potential solution is to define time frames in the consent, to specify in the agreement how long (e.g., 5 years or 10 years or may be lifetime–in such cases, copyright could be assigned to the organization continuing to use the data for life to reduce complexity, following standard copyright transfer procedures as outlined in the U.S. Copyright Act United States Copyright Office (2023)) their data may be used and what would be the protocol after the timeline ends such as establishing a new agreement for continued use or exercising the right to data deletion. We acknowledge that ensuring long-term reliability and verifying proxy legitimacy is challenging. If organizations use post-mortem data, they must enable revocation requests aligned with the consent timeline. Protocols should enforce deletion and safety through privacy-by-design. Once verified, only periodic checks at consent renewal are required, while deletion requests need one-time certification, not lifelong re-verification.
Another potential challenge is ensuring meaningful informed consent for post-mortem data use, given non-experts’ limited understanding of AI and its future developments. This challenge can be addressed by emphasizing transparency and clarity in informed consent. Data usage must be clearly explained in terms of purpose, scope, and potential impact, using accessible language for non-experts. Additionally, any application of the data to new technologies must align with the original consent, ensuring that post-mortem data is used only in ways the individual (or their proxy) authorized.
Appendix D Popular AI Agents
Category AI Agents
General AI Assistants ChatGPT (OpenAI), Google Gemini, Microsoft Copilot, Perplexity, Claude, Grok, DeepSeek, Siri, Alexa, Poe
Routine Tasks / Productivity duckbill, Blockit AI, manus, Cluely, Flow, Genspark, AGI, Inc., KIMI
Note-Taking granola, fireflies.ai, Otter.ai
Hardware PLAUD.AI, Limitless, bee
Communication Superhuman, Grammarly, QuillBot
Consumer Robotics matic, SKILD AI, fauna robotics, Prosper, swish
Physical Intelligence cobot
Travel wanderlog, Wanderboat, autopilot, mindtrip, Layla
Fashion ALTA, Doji
Consumer Advocacy DoNotPay
Home Tasks ohai.ai
Finance Monarch, Copilot, Rocket Money, Betterment
Reading Speechify
Creative Expression: Image/Video Generation & Editing Adobe, Leonardo.Ai, KREA, runway, Pika, Ideogram, Higgsfield, descript, KlingAI, MINIMAX, invideo, FLORA, captions, OpusClip, VEED
Image/Video Models Sora, DeepMind, Reve, Black Forest Labs
Creative Social Platforms CIVITAI, VIGGLE, weights, Cara
Presentations GAMMA, Canva, Figma
Music + Audio suno, udio, Riffusion
Writing Support Grammarly, QuillBot
Voice IIElevenLabs
Physical + Mental Health: Physical Health Curai Health, WHOOP, ROON, Eight Sleep, OURA
Mental Health Woebot Health, Ash, Finch
Nutrition Alma, Cal AI
Learning + Development: Tutoring photomath, Atypical, Gizmo, Super Teacher, Class Companion
App Creation replit, bolt, Framer, Lovable, same, Firebase Studio
Coding Cursor, Windsurf
Language duolingo, Speak
Connection: Dating SITCH, KEEPER, Ditto AI
Spirituality Bible Chat, Hallow, Co-Star
AI Companions character.ai, Replika, Kindroid
Networking gigi
Series BOARDY.AI
Voice sesame
Digital Mind Delphi
Table 1: List of AI agents based on Popularity and Consumers Interests Topics (2025); Carolan et al. (2025).